OpenStudy (anonymous):

the sum of Radical 75 + Radical 3 is?

OpenStudy (mertsj):

Simplify sqrt75 first

OpenStudy (mertsj):

\[\sqrt{75}=\sqrt{25(3)}\]

OpenStudy (anonymous):

so 5

OpenStudy (mertsj):

\[\sqrt{25(3)}+1\sqrt{3}=5\sqrt{3}+1\sqrt{3}=6\sqrt{3}\]

OpenStudy (anonymous):